### v2

# Parallel Split Learning with Global Sampling

Abstract
The expansion of IoT devices and the demands of deep learning have highlighted significant challenges in distributed deep learning systems. Parallel split learning has emerged as a promising derivative of split learning well suited for distributed learning on resource-constrained devices. However, parallel split learning faces several challenges, such as large effective batch sizes, non-independent and identically distributed data, and the straggler effect. We view these issues as a sampling dilemma and propose to address them by orchestrating a mini-batch sampling process on the server side. We introduce a new method called uniform global sampling to decouple the effective batch size from the number of clients and reduce the mini-batch deviation. To address the straggler effect, we introduce a novel method called Latent Dirichlet Sampling, which generalizes uniform global sampling to balance the trade-off between batch deviation and training time. Our simulations reveal that our proposed methods enhance model accuracy by up to 34.1% in non-independent and identically distributed settings and reduce the training time in the presence of stragglers by up to 62%. In particular, Latent Dirichlet Sampling effectively mitigates the straggler effect without compromising model accuracy or adding significant computational overhead compared to uniform global sampling. Our results demonstrate the potential of our methods to mitigate common challenges in parallel split learning.
Index Terms:
Parallel split learning, distributed deep learning, latent dirichlet sampling, uniform global sampling.
I Introduction
In recent years, deep learning (DL) has undergone transformative growth in various domains [1, 2, 3, 4, 5]. Distributed deep learning (DDL) has emerged as a powerful paradigm in the era of the Internet of Things (IoT) and edge computing, addressing the challenges posed by the proliferation of resource-constrained devices and the exponential growth of data at the network edge [6, 7, 8]. This approach enables the distribution of computationally intensive DL tasks across multiple nodes, including IoT devices and edge servers, to overcome the limitations of centralized learning (CL). Using the collective computational power of interconnected devices, DDL not only enhances the efficiency of model training and inference, but also reduces latency, improves privacy, and optimizes resource utilization.
Federated learning (FL) [9] is a prominent framework in DDL. In federated averaging (FedAvg), clients train model copies locally and periodically aggregate these models by averaging their parameters, facilitating global model convergence [9]. However, the high resource demands of FL restrict its implementation in edge computing and IoT environments, primarily due to substantial computational and communication burdens. Strategies such as client selection [10] and model quantization [11] help mitigate these issues but do not resolve them completely.
Split learning (SL) [12, 13] addresses some limitations of FL by partitioning neural network training between client devices and a central server. Clients conduct initial forward propagation (FP) on local data and send intermediate results to the server, which completes the FP, initiates the backward propagation (BP), and returns gradients to the clients to complete the BP. SL enhances data privacy, reduces computational demands on the client side, and increases the convergence speed compared to FL, making it well-suited for edge and IoT applications [12, 14, 15]. However, it incurs high training and communication latency due to its sequential nature [14].
To address training latency in SL, advanced derivatives of SL like split federated learning (SFL) [16, 17, 18] and parallel split learning (PSL) [19, 20, 21, 22, 23] have been proposed. SFL combines the SL and FL techniques to facilitate parallel model training across multiple devices. However, SFL requires that the server maintains individual model copies for each client and the aggregation of updated client models, which complicates scalability.
Figure 1: A schematic illustration of sampling local batch sizes with fixed (a) and variable (b) local batch sizes. The proposed UGS and LDS methods allow for variable local batch sizes that sum up to a fixed global batch size, which enables fine-grained control over the sampling process. This flexibility is crucial for mitigating the challenges of large effective batch sizes, non-IID data, and the straggler effect in PSL.
PSL enhances parallelism without the need for multiple server model instances or client model aggregation. However, PSL still encounters challenges, including the large effective batch size problem [24], also known as the server-client update imbalance problem [22], where the effective batch size scales with the number of clients. Additionally, devices contribute data of varying distributions and sizes, causing batch distributions to deviate from the overall data distribution, which can lead to poor generalization. This issue of non-IID data is a general problem in DDL, not specific to PSL. Another challenge is the straggler effect [25], where slower devices delay the training process due to the need for synchronization, as the processing time of a single batch is determined by the slowest client, impacting throughput and total training time.
To address the large effective batch size problem and the issue of non-IID data, we propose a new method called uniform global sampling (UGS) as an extension to the PSL framework introduced in [19]. UGS allows for variable local batch sizes that sum up to a fixed global batch size. This decouples the number of clients from the effective batch size. By analysing the deviation probability of global batches [^1] We use the term batch to refer to mini-batches in the context of DL. The difference between global and local batches is explained in Section III. from the overall data distribution, we show that the choice of the local batch sizes is crucial for model convergence. In particular, we show that using fixed local batch sizes can increase the deviation probability of the global batches, especially in the presence of non-IID data. In UGS, the local batch sizes are sampled according to client selection probabilities. We show that if the client selection probabilities are proportional to the dataset sizes, the deviation bound of the global batches is at least as tight as for uniform sampling in central learning. Inspired by the Latent Dirichlet Allocation model [26], we introduce a second novel method named Latent Dirichlet Sampling (LDS) that generalizes UGS to address the straggler effect. LDS introduces a Dirichlet prior distribution over the client selection probabilities to control the concentration of stragglers in global batches. The prior distribution is adjusted based on the delay times of the straggler clients and a trade-off hyperparameter. The client selection probabilities are estimated using the expectation-maximization (EM) algorithm for maximum a posteriori (MAP) estimation [27]. LDS balances the trade-off between batch deviation and training time by sampling local batch sizes according to these estimated client selection probabilities.
An overview of our methods is shown in Figure 1. The main contributions of this paper are listed below:
  • We analyse the deviation probability of global batches in PSL and show why the choice of the local batch size is crucial for the convergence of the model.
  • We introduce the UGS method to address the large effective batch size problem and the non-IID data problem in PSL.
  • Motivated by the straggler effect problem, we introduce LDS to balance the trade-off between batch deviation and training time, ensuring efficient and accurate model training.
  • Our simulation results show that the UGS and LDS methods significantly improve PSL accuracy and convergence speed, especially with non-IID data. LDS effectively mitigates the straggler effect without compromising model accuracy or adding significant computational overhead compared to UGS.
The remainder of this paper is organized as follows. Sections II and III present related works and preliminaries of our work. In Section IV, our novel UGS and LDS methods are introduced after analyzing the mini-batch deviation in PSL. Section V presents the simulation results and Section VI concludes the paper with final remarks.
II Related Work
Initial versions of PSL faced the client decoupling problem [24] (also known as client detachment problem [22]), where the information flow between clients is decoupled during the BP due to independent execution. Jeon and Kim [19] use gradient averaging over all clients to address the client decoupling problem. They chose the local batch sizes fixed and proportional to the dataset size of each client. Lyu et al. [28] propose averaging the clients’ local loss function outputs instead of their local gradients. This approach aligns with that of [19] due to the linearity of the gradient operator, but enhances privacy as clients can compute the local loss functions without sharing labels. Both show consistent convergence and improved accuracy in both IID and non-IID settings but do not address the large effective batch size problem.
Pal et al. [24] propose gradient averaging over a subset of the clients. Opposed to [19] and [28], the authors observe that the model may not converge if gradients are averaged over all clients. To address the large effective batch size problem, the authors propose scaling the learning rate on the server side while keeping the local batch size fixed and identical for all clients. Following the line of work from [23], Oh et al. [22] propose using smashed data augmentation, local loss function design and FedAvg for the client models to overcome the large effective batch size problem and client model decoupling.
Pal et al. [24] and Oh et al. [22] provide promising solutions for the large effective batch size problem, but assume that the effective batch size scales with the number of clients. However, it has been shown that increasing the batch size leads to diminishing returns for gradient estimation [29] and that smaller batch sizes [30] can be more beneficial for model generalization. Furthermore, the empirical results are reported only for IID settings. They do not address the non-IID data problem, which is a significant challenge in PSL.
Cai and Wei [31] propose an incentive mechanism to handle non-IID data in PSL, which encourages clients with larger and more evenly distributed datasets to participate in training. The bargaining game approach for personalized SL [32] balances energy consumption and privacy in non-IID settings. However, these methods may face scalability issues and may not fully address the heterogeneity of data distributions across diverse clients.
Other efforts to improve PSL focus on reducing training and communication latency [20, 21, 33]. Methods include last-layer gradient aggregation, resource optimization, client clustering, and cut-layer selection. However, these methods do not address the straggler effect, which remains a significant challenge in PSL. A notable related work is the BOA algorithm [34], which dynamically adjusts the batch size of slow clients in distributed learning systems to reduce the straggler effect.
Contrary to other works, we view the large effective batch size problem, the non-IID data problem, and the straggler effect as a sampling dilemma. By orchestrating the batch sampling process on the server side, we allow variable local batch sizes and decouple the number of clients from the effective batch size. This enables global sampling methods to account for non-IID data by reducing the deviation between the global batches and the overall data distribution and mitigating the straggler effect by controlling the concentration of stragglers in global batches.
III Preliminaries
We consider the PSL setting from [19] where a server and $K\in\mathbb{N}$ clients train a neural network model with parameter vector $\boldsymbol{w}$ collaboratively using the PSL framework. The set of clients is represented by $\mathcal{K}=\{1,2,\ldots,K\}$. We focus on supervised classification tasks, where the dataset $\mathcal{D}_{k}$ of each client $k\in\mathcal{K}$ consists of input-label pairs $\mathcal{D}_{k}=\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{D_{k}}$, where $D_{k}\in\mathbb{N}$ is the dataset size, $\boldsymbol{x}_{k,i}$ is the input data, and $y_{k,i}\in\mathcal{M}$ is the ground truth class label of $\boldsymbol{x}_{k,i}$. Here, =$\mathcal{M}=\{1,2,\dots,M\}$ represents all classes and $M\in\mathbb{N}$ is the number of classes. The overall dataset $\mathcal{D}_{0}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$ is the union of all clients’ datasets, with a total size $D_{0}=\sum_{k=1}^{K}D_{k}$. We refer to the distribution of the data points in $\mathcal{D}_{0}$ as the empirical distribution. We define the class counting function $c:\mathbb{N}^{d}\times\mathbb{N}\to\mathbb{N}$ with $d\in\mathbb{N}$ as
$c(\boldsymbol{y},m)=\sum_{i=1}^{d}\delta(y_{i},m)$
$\text{where}\quad\delta(i,j)=\begin{cases}1&\text{if }i=j\\ 0&\text{otherwise}\end{cases}$
The class distribution of a client $k$ with class label vector $\boldsymbol{y}_{k}\in\mathcal{M}^{D_{k}}$ is denoted as $\boldsymbol{\beta}_{k}\in\mathbb{R}^{M}$ with ${\boldsymbol{\beta}_{k}=\frac{1}{D_{k}}(c(\boldsymbol{y}_{k},1),c(\boldsymbol{y}_{k},2),\dots,c(\boldsymbol{y}_{k},M))^{\top}}$. The overall class distribution of $\mathcal{D}_{0}$ with class label vector $\boldsymbol{y}_{0}\in\mathcal{M}^{D_{0}}$ is denoted as $\boldsymbol{\beta}_{0}\in\mathbb{R}^{M}$ according to the same definition.
Every epoch consists of $T\in\mathbb{N}$ optimization steps. The set of optimization steps is denoted as $\mathcal{T}=\{1,2,\ldots,T\}$. We use the subscript $k$ for clients and the superscript $t$ for optimization steps. If the distinction is unnecessary, we omit the indices for brevity. Each client has a local batch size $B_{k}\in\mathbb{N}$. The global batch size $B=\sum_{k=1}^{K}B_{k}$ is the total number of data points that arrive at the server in each optimization step. Accordingly, we differentiate between local batches $\mathcal{B}^{(t)}_{k}\subseteq\mathcal{D}_{k}$ and the global batch $\mathcal{B}^{(t)}=\bigcup_{k=1}^{K}\mathcal{B}^{(t)}_{k}$ in each optimization step.
PSL mirrors SL by splitting the model parameters into a client-side lower segment $\boldsymbol{w}_{c}$ and a server-side upper segment $\boldsymbol{w}_{s}$ with $\boldsymbol{w}=(\boldsymbol{w}_{s},\boldsymbol{w}_{c})^{\top}$. The first layer on the server side is referred to as the cut layer. Before training, the server sends $\boldsymbol{w}_{c}$ to the clients, each storing a copy $\boldsymbol{w}_{c,k}$ with identical initial values. The function $f(\boldsymbol{x};\boldsymbol{w})$ maps $\boldsymbol{x}$ to a predicted value according to the given model parameters $\boldsymbol{w}$. Hence, the global model is given as $f(\boldsymbol{x};\boldsymbol{w})=f(f(\boldsymbol{x};\boldsymbol{w}_{c});\boldsymbol{w}_{s})$. When the argument of $f$ is a dataset or a subset of it, e.g., a batch $\mathcal{B}$ with batch size $B$, it operates element-wise on the input data and is defined as:
$f(\mathcal{B};\boldsymbol{w})=\{(f(\boldsymbol{x}_{i};\boldsymbol{w}),y_{i})\}_{i=1}^{B}$
The goal is to find the optimal model parameters ${\boldsymbol{w}^{*}=(\boldsymbol{w}_{s}^{*},\boldsymbol{w}_{c}^{*})^{\top}}$ that minimize the global loss function $L(\mathcal{D}_{0};\boldsymbol{w})$:
$\boldsymbol{w}^{*}$ $=\operatorname*{arg\,min}_{\boldsymbol{w}}L(\mathcal{D}_{0};\boldsymbol{w})$
$=\operatorname*{arg\,min}_{\boldsymbol{w}}\frac{1}{D_{0}}\sum_{k=1}^{K}\sum_{i=1}^{D_{k}}L((\boldsymbol{x}_{k,i},y_{k,i});\boldsymbol{w})$
$=\operatorname*{arg\,min}_{\boldsymbol{w}}\frac{1}{D_{0}}\sum_{k=1}^{K}\sum_{i=1}^{D_{k}}L((f(\boldsymbol{x}_{k,i};\boldsymbol{w}_{c}),y_{k,i});\boldsymbol{w}_{s})$
The mini-batch stochastic gradient descent (SGD) algorithm or one of its variants are typically used to optimize the model parameters, alternating between FP and BP. Each optimization step $t\in\mathcal{T}$ involves the following substeps.
1) Parallel Client-Side FP
Each client $k$ samples a local batch $\mathcal{B}^{(t)}_{k}\subseteq\mathcal{D}_{k}$ with a local batch size of $B_{k}$ uniformly at random without replacement and sends $f(\mathcal{B}^{(t)}_{k};\boldsymbol{w}_{c})$ to the server.
2) Server-Side FP
Upon receiving all intermediate results from the clients, the server concatenates them to obtain the global batch ${\mathcal{B}^{(t)}=\bigcup_{k=1}^{K}f(\mathcal{B}^{(t)}_{k};\boldsymbol{w}_{c})}$. The server then computes the final prediction-label pairs $f(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$ to calculate the average loss $L(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$.
3) Server-Side BP
The server computes the gradient $\boldsymbol{g}^{(t)}_{\boldsymbol{w}_{s}}=\nabla_{\boldsymbol{w}_{s}}L(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$. Using the learning rate $\eta\in\mathbb{R}$, the server updates its model parameters to $\boldsymbol{w}^{(t+1)}_{s}=\boldsymbol{w}^{(t)}_{s}-\eta\boldsymbol{g}^{(t)}_{\boldsymbol{w}_{s}}$. The server then broadcasts the gradient of the cut layer to all clients.
4) Parallel Client-Side BP
Upon receiving the cut layer gradient, each client $k$ computes the gradient $\boldsymbol{g}^{(t)}_{\boldsymbol{w}_{c,k}}=\nabla_{\boldsymbol{w}_{c}}L(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$ and sends it back to the server.
5) Server-Side Gradient Averaging
Upon receiving all gradients from the clients, the server computes the average client gradient $\boldsymbol{\overline{g}}^{(t)}_{\boldsymbol{w}_{c}}=\sum_{k=1}^{K}\frac{D_{k}}{D_{0}}\boldsymbol{g}^{(t)}_{\boldsymbol{w}_{c,k}}$ and broadcasts it to all clients.
6) Parallel Client-Side Model Update
Each client $k$ uses $\boldsymbol{\overline{g}}^{(t)}_{\boldsymbol{w}_{c}}$ to update its model parameters $\boldsymbol{w}^{(t+1)}_{c,k}=\boldsymbol{w}^{(t)}_{c,k}-\eta\boldsymbol{\overline{g}}^{(t)}_{\boldsymbol{w}_{c}}$, using the same learning rate $\eta$ as the server.
An epoch is concluded after $T$ optimization steps. $T$ is determined by the number of non-empty global batches that can be formed from the clients’ datasets until depletion. The model is trained for a fixed number of epochs $E\in\mathbb{N}$. Note that since the client models are initialized with the same parameters $\boldsymbol{w}_{c}$ and are always updated with the same gradients $\boldsymbol{\overline{g}}^{(t)}_{\boldsymbol{w}_{c}}$, they remain identical throughout the training process.
We make the following simplifying assumptions throughout this work:
Availability of Client Classes and Dataset Sizes: The server has access to the class counts and dataset sizes of all clients. It is important to note that in all SL-based frameworks, the class labels are shared during training and can be easily collected by the server. Therefor, providing the class counts or dataset sizes to the server does not introduce additional privacy concerns. The class distributions can be inferred from the class counts and dataset sizes.
Multinomial Approximation of Uniform Sampling: We model the distribution of local batches that are sampled uniformly at random without replacement as a multinomial distribution. The correct model would be a hypergeometric distribution. However, assuming that usually $B_{k}\ll D_{k}$, the multinomial distribution closely approximates the hypergeometric distribution.
Stable Per-Epoch Delay Times: We assume that the delay times of the straggler clients do not change significantly during an epoch. The server keeps an updated estimate of the delay times $\omega_{k}\in\mathbb{R}_{\geq 0}$ for each client $k$. We define the delay time of a client as the difference between the processing time of the client and the processing time of the fastest client. The delay times are always relative to the fastest client, i.e., the fastest client $k^{\prime}$ has a delay time of $\omega_{k^{\prime}}=0$. The server may initially set the delay times to zero, if no prior knowledge is available
A summary of notations used in this work can be found in Table I.
TABLE I: Summary of Notations.
Notation Description
$\mathcal{K}$ Index set for the $K$ clients.
$\mathcal{D}_{k}$ Dataset of client $k$. $k=0$ for overall dataset.
$D_{k}$ Dataset size of dataset $\mathcal{D}_{k}$. $k=0$ for overall dataset.
$\mathcal{M}$ Set of all $M$ unique labels.
$c$ Class counting function.
$\beta_{k}$ Class distribution of client $k$. $k=0$ for overall class distribution.
$\omega_{k}$ Delay time of client $k$ relative to the fastest one.
$L(\mathcal{D}_{0};\boldsymbol{w})$ Global loss function.
$f$ Model mapping function.
$\boldsymbol{w}$ Global model parameter vector.
$\boldsymbol{w}_{s}$/$\boldsymbol{w}_{c}$ Server/client model parameter vectors.
$\mathcal{T}$ Set of optimization steps. $T$ in total.
$\mathcal{B}_{k}^{(t)}$ Local batch from client $k$ at optimization step $t$.
$B_{k}$ Local batch size of client $k$.
$\mathcal{B}^{(t)}$ Global batch at optimization step $t$
$B$ Global batch size. $B=\sum_{k=1}^{K}B_{k}$.
$\boldsymbol{g}_{\boldsymbol{w}}^{(t)}$ Gradient vector of $L(\mathcal{B}^{(t)};\boldsymbol{w})$.
$\eta$ Learning rate for server and clients.
$E$ Number of epochs/training rounds.
IV Parallel Split Learning with Global Sampling
A good sampling method should ensure that the expected distribution of the global batches aligns with the empirical distribution. If the global batches frequently deviate from the empirical distribution, the trained model may become biased and generalize poorly. We measure the quality of a sampling method by the deviation of the generated global batches from the empirical distribution, referred to as batch-deviation. By analyzing how the choice of local batch sizes affects the batch deviation, we determine the UGS method that is bounded at least as tight as to central uniform sampling in terms of batch deviation and decouples the effective batch size from the number of clients. All proofs are provided in Appendix B.
IV-A Deviation Analysis
Assume we are given the complete dataset $\mathcal{D}_{0}$ and train our model using central learning. In this scenario, the model is typically trained using batches $\mathcal{B}$ of size $B$ that are sampled uniformly at random. Although uniform sampling does not guarantee minimal batch deviation, it is common due to its simplicity and efficiency. To simplify our analysis, we consider the overall class distribution $\boldsymbol{\beta}_{0}$ instead of the empirical distribution, assuming that the class distribution acts as a proxy for the empirical distribution. We define the batch deviation as the $L_{1}$ distance between the class distribution of a batch $\mathcal{B}$ with class label vector $\boldsymbol{y}$ and $\boldsymbol{\beta}_{0}$:
$d(\mathcal{B},\boldsymbol{\beta}_{0})=\sum_{m=1}^{M}\left|\frac{c(\boldsymbol{y},m)}{B}-\beta_{0,m}\right|$ (1)
Let ${Y=(Y_{1},Y_{2},\dots,Y_{M})^{\top}}$ be a random vector representing the class counts in a batch $\mathcal{B}$ sampled from $\mathcal{D}_{0}$ uniformly at random with $Y\sim\text{Multinomial}(B,\boldsymbol{\beta}_{0})$. Using the Chebychev inequality, we can derive an upper bound on the probability of $\mathcal{B}$ deviating from $\boldsymbol{\beta}_{0}$ by at least $\epsilon$:
Lemma 1.
For all $\epsilon>0$ and $m\in\mathcal{M}$ it holds that
$P\left(\left|\frac{Y_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}.$
Our goal is to determine a sampling method for PSL with a deviation bound at least as tight as the one from central uniform sampling. In the regular PSL framework, a local batch size $B_{k}$ is assigned to each client $k$ and does not change during training. The class distribution of each local batch $\mathcal{B}^{\prime}_{k}$ is modelled as a multinomial distribution with parameters $B_{k}$ and $\boldsymbol{\beta}_{k}$. Let $Y^{\prime}=(Y^{\prime}_{1},Y^{\prime}_{2},\dots,Y^{\prime}_{M})^{\top}$ be a random vector representing the class counts in the corresponding global batch $\mathcal{B}^{\prime}$. Since the $\mathcal{B}^{\prime}_{k}$ are sampled independently, $Y^{\prime}$ is distributed according to the sum of the multinomials:
$Y^{\prime}\sim\sum_{k=1}^{K}\text{Multinomial}(B_{k},\boldsymbol{\beta}_{k}).$
Using the Markov inequality, we can derive an upper bound on the probability of $\mathcal{B}^{\prime}$ deviating from $\boldsymbol{\beta}_{0}$ by at least $\epsilon$ and relate it to $\mathcal{B}$, i.e., batches generated by central uniform sampling:
Lemma 2.
For all $\epsilon>0$ and $m\in\mathcal{M}$ it holds that
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y^{\prime}_{m})+(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}])^{2}}{B^{2}\epsilon^{2}}.$
The upper bound loosens when the bias term ${(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}])^{2}}$ increases. Since $\mathbb{E}[Y^{\prime}_{m}]=\sum_{k=1}^{K}B_{k}\beta_{k,m}$ and $\mathbb{E}[Y_{m}]=B\beta_{0,m}$, this implies that even small differences in the $B_{k}$ or $\beta_{k,m}$ can lead to a significant deviation in the global batches.
Theorem 1.
If $B_{k}=B\frac{D_{k}}{D_{0}}$ or $\beta_{k,m}=\beta_{0,m}$ for all $k\in\mathcal{K}$ and $m\in\mathcal{M}$, then it holds that
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}$
for all $m\in\mathcal{M}$.
Theorem 1 shows that choosing the $B_{k}$ proportionally to the local dataset sizes ensures that the upper bound is at least as tight as the one from central uniform sampling in Lemma 1. Only if the client datasets are IID ($\beta_{k,m}=\beta_{0,m}$), any viable choice of local batch sizes $B_{k}$ ensures the tightness of the deviation bound. Therefore, choosing the same local batch size $B_{k}=B^{\prime}$ for all clients $k$ as in [24] can lead to an increased batch deviation when the client datasets are non-IID. On the other hand, if $B\frac{D_{k}}{D_{0}}\notin\mathbb{N}$, the local batch sizes must be converted to integers. This can increase the deviation of the global batches in a non-IID setting, since the premise of Theorem 1 may not hold anymore throughout the training process.
IV-B Uniform Global Sampling (UGS)
We propose the UGS method, which ensures that Theorem 1 holds independent of local dataset sizes or local data distributions. The complete UGS method is outlined in Algorithm 1. In UGS, the local batch sizes are variable and determined by the server for each training step separately. It is not required to fix the local batch sizes right from the start. The server only requires a global batch size $B$ to be specified.
UGS is invoked before each epoch of PSL. Each client $k$ receives a local batch size vector $(B_{k}^{(1)},B_{k}^{(2)},\dots,B_{k}^{(T)})^{\top}$ from the server such that ${\sum_{k=1}^{K}B_{k}^{(t)}=B}$ for all $t\in\mathcal{T}$. The $B_{k}^{(t)}$ specify the local batch size of client $k$ at training step $t$. If $B_{k}^{(t)}=0$, the client $k$ does not contribute to the global batch at training step $t$. UGS uses the following generative process to sample the global batches $\mathcal{B}^{(t)}$ in each training step $t$:
  1. 1.
    Choose $\boldsymbol{\pi}=\left(\frac{D_{1}}{D},\frac{D_{2}}{D},\dots,\frac{D_{K}}{D}\right)^{\top}$
  2. 2.
    For each $i\in\{1,2,\ldots,B\}$:
    1. (a)
      Choose a client $z_{i}^{(t)}\sim\text{Categorical}(\boldsymbol{\pi})$
    2. (b)
      Choose a class $y_{i}^{(t)}\sim\text{Categorical}(\boldsymbol{\beta}_{z_{i}^{(t)}})$
The key idea is to first sample a client $z_{i}^{(t)}\in\mathcal{K}$ according to a categorical distribution with parameters $\boldsymbol{\pi}$ and then sample a class $y_{i}^{(t)}\in\mathcal{M}$ from the chosen client. Since $\sum_{k=1}^{K}\pi_{k}=1$, UGS corresponds to sampling from a mixture of categorical distributions with the mixing probabilities $\boldsymbol{\pi}$. In our context, $\boldsymbol{\pi}$ can be interpreted as the client selection probabilities. Note that the categorical distribution is used iteratively instead of sampling in one step from the multinomial distribution. This is because the clients actually sample uniformly without replacement. If a client $k$ has no data left, the server needs to update the client selection probabilities $\boldsymbol{\pi}$ by setting $\pi_{k}=0$ and renormalizing the remaining probabilities before continuing the sampling process.
$\boldsymbol{z}^{(t)}=(z_{1}^{(t)},{z}_{2}^{(t)},\dots,{z}_{B}^{(t)})^{\top}$ represent the clients from which the $y_{i}^{(t)}$ are sampled. This allows the sampling process to be split between the server and the clients, where the server executes 2a) and the clients execute 2b). Since steps 2a) and 2b) are independent, we can postpone step 2b) and pre-compute the $B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$. Accordingly, we count the occurrences of each client $k$ in $\boldsymbol{z}^{(t)}$ with $B_{k}^{(t)}=c(\boldsymbol{z}^{(t)},k)$ right after step 2a). The server then sends the $B_{k}^{(t)}$ to the clients.
Once training starts, the clients can execute step 2b) on demand, i.e., sampling local batches of size $B_{k}^{(t)}$. This equals substituting $B_{k}$ with $B_{k}^{(t)}$ in the first step of PSL in Section III. The rest of the training process remains unchanged. Consequently, the effective batch size $B$ becomes a hyperparameter that is no longer dependent on the number of clients or their respective dataset sizes, addressing the large effective batch size problem.
Corollary 1.
Let $\boldsymbol{\pi}=\left(\frac{D_{1}}{D_{0}},\frac{D_{2}}{D_{0}},\dots,\frac{D_{K}}{D_{0}}\right)^{\top}$. If
$Y^{\prime}_{m}\sim B\sum_{k=1}^{K}\pi_{k}\textnormal{Categorical}(\boldsymbol{\beta}_{k})$
then
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}$
for all $m\in\mathcal{M}$.
Corollary 1 shows that Theorem 1 holds for the global batches generated by this process. Therefor, UGS ensures that the batch deviation bound is as tight as the one from central uniform sampling even in the presence of non-IID data distributions.
Algorithm 1 Uniform Global Sampling
1:$(D_{1},\dots,D_{K})^{\top},B$
2:$B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$
3:$D\leftarrow\sum_{k=1}^{K}D_{k}$
4:$T\leftarrow\left\lceil D/B\right\rceil$
5:$B_{k}^{(t)}\leftarrow 0$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$
6:
7:$\boldsymbol{\pi}\leftarrow(\frac{D_{1}}{D},\frac{D_{2}}{D},\dots,\frac{D_{K}}{D})^{\top}$
8:
9:for each training step $t=1$ to $T$ do
10:     for $i=1$ to $B$ do
11:         $z_{i}\sim\text{Categorical}(\boldsymbol{\pi})$
12:         $B_{z_{i}}^{(t)}\leftarrow B_{z_{i}}^{(t)}+1$
13:         if $\sum_{t^{\prime}=1}^{t}B_{z_{i}}^{(t^{\prime})}=D_{z_{i}}$ then
14:              $\boldsymbol{\pi}_{z_{i}}\leftarrow 0$
15:              $\boldsymbol{\pi}\leftarrow\boldsymbol{\pi}/\sum_{k=1}^{K}\boldsymbol{\pi}_{k}$
16:         end if
17:     end for
18:end for
19:return $B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$
IV-C Straggler Mitigation
While UGS addresses the issue of large effective batch sizes and non-IID data distributions, it does not consider the presence of stragglers. Intuitively, by increasing the client selection probability $\pi_{k}^{\prime}$ of a straggler client $k^{\prime}$, we can reduce the number of global batches to which the straggler client contributes. This reduces the total training time, as the straggler clients’ datasets deplete faster and they appear less frequently in global batches.
However, adjusting the $\boldsymbol{\pi}$ can increase the batch deviation, as Corollary 1 would no longer hold. The choice of $\boldsymbol{\pi}$ is thus crucial to balance the trade-off between the batch deviation and the concentration of stragglers in the global batches. To address this issue, we treat $\boldsymbol{\pi}$ as a parameter to be optimized with respect to the desired trade-off. Inspired by the Latent Dirichlet Allocation model [26], we introduce a Dirichlet prior distribution over $\boldsymbol{\pi}$ with concentration parameter $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ and $\alpha_{k}>0$ and treat $\boldsymbol{z}$ as a latent variable. As a conjugate to the categorical distribution, the Dirichlet distribution is a suitable choice for the prior distribution. The choice of $\boldsymbol{\alpha}$ is crucial to balance the trade-off between batch deviation and total training time. For now, we assume that $\boldsymbol{\alpha}$ is given and will discuss its initialization in the next subsection. The prior distribution over $\boldsymbol{\pi}$ is defined as:
$P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})=\frac{1}{\textnormal{B}(\boldsymbol{\alpha})}\prod_{k=1}^{K}\pi_{k}^{\alpha_{k}-1}$
where $\textnormal{B}(\boldsymbol{\alpha})$ is the Beta function. Furthermore we have:
$p(z\mid\boldsymbol{\pi})=\pi_{z}\quad\text{and}\quad p(y\mid z,\boldsymbol{\beta})=\beta_{z,y}$
The goal is to estimate the mixture proportions $\boldsymbol{\pi}^{*}$ that maximize the posterior distribution $P(\boldsymbol{\pi}\mid\boldsymbol{y},\boldsymbol{\alpha},\boldsymbol{\beta})$, so that the global batches that are generated using $\boldsymbol{\pi}^{*}$ are less likely to deviate from $\boldsymbol{y}$ while respecting the prior beliefs given by $\boldsymbol{\alpha}$. Here $\boldsymbol{y}\in\mathcal{M}^{N}$ with sample size $N\in\mathbb{N}$ is the observed and incomplete data we want to fit our model to. For an accurate estimation the class distribution of $\boldsymbol{y}$ should match the overall class distribution as closely as possible. This can be formulated as a MAP estimation problem:
$\boldsymbol{\pi}^{*}$ $=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\>\>P(\boldsymbol{\pi}\mid\boldsymbol{y},\boldsymbol{\alpha},\boldsymbol{\beta})$
$=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\>\>P(\boldsymbol{y}\mid\boldsymbol{\pi},\boldsymbol{\beta})\cdot P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$
$=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\>\>\ln P(\boldsymbol{y}\mid\boldsymbol{\pi},\boldsymbol{\beta})+\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$
$=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\>\>\ln\Big{\{}\sum_{\boldsymbol{z}\in\mathcal{K}^{N}}P(\boldsymbol{y},\boldsymbol{z}\mid\boldsymbol{\pi},\boldsymbol{\beta})\Big{\}}+\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$ (2)
Due to the sum inside the logarithm in Equation 2, a closed-form solution for the MAP estimate is not available. Instead, we use the EM algorithm [27] to iteratively estimate the mixture proportions $\boldsymbol{\pi}^{*}$.
The EM algorithm is a two-step iterative optimization algorithm that alternates between the expectation (E) step and the maximization (M) step to seek the maximum likelihood estimate of a marginal likelihood function. We use the MAP variant of the EM algorithm. It starts with an initial parameter value $\boldsymbol{\pi}^{\text{old}}\sim P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$ and iteratively refines the estimate until convergence. The E-step computes the expected value of the log likelihood $P(\boldsymbol{y},\boldsymbol{z}\mid\boldsymbol{\pi},\boldsymbol{\beta})$ w.r.t. $\boldsymbol{\pi}^{\text{old}}$ and is given by:
$\mathcal{Q}(\boldsymbol{\pi},\boldsymbol{\pi}^{\text{old}})$ $=\mathbb{E}_{\boldsymbol{z}}\Big{[}\ln P(\boldsymbol{y},\boldsymbol{z}\mid\boldsymbol{\pi},\boldsymbol{\beta})\Big{]}$
$=\sum_{\boldsymbol{z}\in\mathcal{K}^{N}}P(\boldsymbol{z}\mid\boldsymbol{y},\boldsymbol{\pi}^{\text{old}},\boldsymbol{\beta})\ln P(\boldsymbol{y},\boldsymbol{z}\mid\boldsymbol{\pi},\boldsymbol{\beta})$
$=\sum_{\boldsymbol{z}\in\mathcal{K}^{N}}P(\boldsymbol{z}\mid\boldsymbol{y},\boldsymbol{\pi}^{\text{old}},\boldsymbol{\beta})\ln\big{\{}P(\boldsymbol{y}\mid\boldsymbol{z},\boldsymbol{\beta})P(\boldsymbol{z}\mid\boldsymbol{\pi})\big{\}}$
$=\sum_{n=1}^{N}\sum_{k=1}^{K}\gamma_{nk}\big{\{}\ln\pi_{k}+\ln\beta_{k,y_{n}}\big{\}}.$ (3)
Note that we use the expected value of the posterior distribution $P(\boldsymbol{z}\mid\boldsymbol{y},\boldsymbol{\pi}^{\text{old}},\boldsymbol{\beta})$ in Equation 3 since the latent variable $\boldsymbol{z}$ is not observed. The responsibilities $\gamma_{nk}$ are given by:
$\gamma_{nk}=\frac{\pi_{k}^{\text{old}}\beta_{k,y_{n}}}{\sum_{k^{\prime}\in\mathcal{K}}\pi_{k^{\prime}}^{\text{old}}\beta_{k^{\prime},y_{n}}}.$
In the M-step, $\mathcal{Q}$ is maximized with respect to $\boldsymbol{\pi}$. Adding the log prior $\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$ to the objective function is the only modification required to use the EM algorithm for MAP estimation [27]. The M-step is given by:
$\boldsymbol{\pi}^{\text{new}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\mathcal{Q}(\boldsymbol{\pi},\boldsymbol{\pi}^{\text{old}})+\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha}).$ (4)
Solving the maximization problem yields the following update rule:
Proposition 1.
The M-step for Equation 4 has a closed-form solution and is given by:
$\pi_{k}^{\textnormal{new}}=\frac{N_{k}+\alpha_{k}-1}{N+\alpha_{0}-K}$
where $N_{k}=\sum_{n=1}^{N}\gamma_{nk}$ and $\alpha_{0}=\sum_{k=1}^{K}\alpha_{k}$.
Afterwards, $\boldsymbol{\pi}^{\text{old}}$ is updated to $\boldsymbol{\pi}^{\text{new}}$ and the E- and M-steps are iteratively repeated until $||\boldsymbol{\pi}^{\text{new}}-\boldsymbol{\pi}^{\text{old}}||_{2}<\tau$, where $\tau>0$ is the convergence threshold hyperparameter. The complete EM algorithm for the MAP estimation is outlined in Algorithm 2.
Note, that since $y_{n}=y_{n^{\prime}}\implies\gamma_{nk}=\gamma_{n^{\prime}k}$ for $n,n^{\prime}\in\{1,2,\dots,N\}$, the responsibilities $\gamma_{nk}$ can be precomputed for each class $m\in\mathcal{M}$. We define $\hat{\boldsymbol{\gamma}}_{k}\in\mathbb{R}^{M}$ as the classwise responsibilities for client $k$ with
$\hat{\gamma}_{k,m}=\frac{\pi_{k}^{\text{old}}\beta_{k,m}}{\sum_{k^{\prime}\in\mathcal{K}}\pi_{k^{\prime}}^{\text{old}}\beta_{k^{\prime},m}}$ (5)
and $\boldsymbol{\nu}\in\mathbb{N}^{M}$ as the class counts in $\boldsymbol{y}$ with ${\nu_{m}=(c(\boldsymbol{y},1),c(\boldsymbol{y},2),\dots,c(\boldsymbol{y},M))^{\top}}$.
This allows us to reformulate the term $N_{k}$ in the M-step as $N_{k}=\boldsymbol{\nu}^{\top}\hat{\boldsymbol{\gamma}}_{k}$. The class counts $\boldsymbol{\nu}$ need to be computed only once at the beginning of the EM algorithm and $\hat{\boldsymbol{\gamma}}_{k}$ involves the computation of $M\cdot K$ responsibilities per iteration. In the original formulation, $N\cdot K$ responsibilities need to be computed per iteration. Since usually $M\ll N$, this reformulation can lead to a significant speedup. This means that we can take advantage of the complete dataset $\mathcal{D}_{0}$ by choosing $\boldsymbol{y}=\boldsymbol{y}_{0}$ for a more accurate estimation of $\boldsymbol{{\pi}}^{*}$ without additional computational cost compared to using an $N$-sized sample (except for the computation of $\boldsymbol{\nu}$, which is negligible).
Algorithm 2 EM Algorithm for MAP Estimation
1:$\boldsymbol{y},\boldsymbol{\pi},\boldsymbol{\beta},\boldsymbol{\alpha},\tau$
2:$\boldsymbol{\pi}^{*}$
3:$\boldsymbol{\pi}^{\text{old}}\leftarrow\boldsymbol{\pi}\cdot\infty$
4:$\boldsymbol{\pi}^{\text{new}}\leftarrow\boldsymbol{\pi}$
5:$N\leftarrow|\boldsymbol{y}|$
6:$\alpha_{0}\leftarrow\sum_{k=1}^{K}\alpha_{k}$
7:$\boldsymbol{\nu}\leftarrow(c(\boldsymbol{y},1),c(\boldsymbol{y},2),\dots,c(\boldsymbol{y},M))^{\top}$
8:
9:while $||\boldsymbol{\pi}^{\text{new}}-\boldsymbol{\pi}^{\text{old}}||_{2}\geq\tau$ do
10:     $\boldsymbol{\pi}^{\text{old}}\leftarrow\boldsymbol{\pi}^{\text{new}}$
11:     for $k=1$ to $K$ do
12:         Compute $\hat{\boldsymbol{\gamma}}_{k}$ using (5)
13:         $N_{k}\leftarrow\boldsymbol{\nu}^{\top}\hat{\boldsymbol{\gamma}}_{k}$triangleright E-Step
14:         $\boldsymbol{\pi}_{k}^{\text{new}}\leftarrow\frac{N_{k}+\alpha_{k}-1}{N+\alpha_{0}-K}$triangleright M-Step
15:     end for
16:end while
17:return $\boldsymbol{\pi}^{\text{new}}$
IV-D Initialization of Concentration Parameters
The EM algorithm is sensitive to the initialization of the parameters and may converge to different local maxima depending on the initialization. We initialize $\boldsymbol{\alpha}$ in two steps. First, we set $\alpha_{k}=\frac{D_{k}}{D}\cdot N$. This initial step is independent of the stragglers and provides a good starting point for optimization. The expected value for $\pi_{k}$ with this initialization is $\frac{D_{k}}{D}$, which aligns with Corollary 1 and ensures that the $\pi_{k}$ are most likely proportional to the dataset sizes. Our heuristic initialization keeps the $\alpha_{k}$ values in a range where they neither dominate the $N_{k}$ in Proposition 1 nor are they insignificant, ensuring $N_{k}\leq\alpha_{k}$.
In the second step, we adjust the $\alpha_{k}$ for each client $k$ to account for the delay times given a hyperparameter $\Delta\geq 0$:
$\alpha_{k}\leftarrow\alpha_{k}\cdot\exp\{\Delta\cdot\frac{(\omega_{k}-\bar{\omega})}{s_{\boldsymbol{\omega}}}\}$
$\text{where}\quad\bar{\omega}=\frac{1}{K}\sum_{k=1}^{K}\omega_{k}\quad\text{and}\quad s_{\boldsymbol{\omega}}=\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}(\omega_{k}-\bar{\omega})^{2}}$
In this adjustment, the z-score $\frac{\omega_{k}-\bar{\boldsymbol{\omega}}}{s_{\boldsymbol{\omega}}}$ standardizes the delay times $\omega_{k}$ by measuring how many standard deviations each delay time is from the mean delay time $\bar{\omega}$. By incorporating the z-score into the exponential function, we achieve a smooth adjustment that either increases or decreases $\alpha_{k}$ based on the relative delay time of each client. $\Delta$ is the single hyperparameter that controls the trade-off between the batch deviation and the total training time. It changes the unit of measurement of the z-score in terms of standard deviations. Higher values of $\Delta$ lead to a more aggressive adjustment of the $\alpha_{k}$ and can reduce the total training time, but may also increase the batch deviation and vice versa.
IV-E Latent Dirichlet Sampling (LDS)
Figure 2: Number of EM iterations in LDS for different parameter choices. $N$ is given relative to $D_{0}$. (a) $p_{s}=0.1$ (b) $p_{s}=0.2$
We introduce a new method, which we call Latent Dirichlet Sampling (LDS), that balances the trade-off between batch deviation and total training time. LDS is detailed in Algorithm 3. The procedure is almost identical to UGS, but differs in a few aspects.
First, the client selection probabilities $\boldsymbol{\pi}$ are estimated using the EM algorithm for MAP estimation as described in the previous section. Therefor, the trade-off hyperparameter $\Delta$ needs to be provided. The concentration parameter $\boldsymbol{\alpha}$ is initialized as described in the previous section. Note that for $\Delta=0$, the delay times have no impact on $\alpha_{k}$ and the initialization is equivalent to the first step. In this case, the EM algorithm will almost surely converge to the same $\boldsymbol{\pi}$ as in UGS. Thus, UGS can be seen as a special case of LDS with $\Delta=0$.
Second, $\boldsymbol{\pi}$ can not be simply renormalized after the dataset of a client $k$ depletes. Instead, the component $k$ is removed from the model and the remaining $\boldsymbol{\pi}$ are re-estimated to ensure that $\boldsymbol{\pi}$ remains a maximum of the posterior distribution. The choice for $\boldsymbol{\pi}$ at the beginning of the re-estimation is a crucial factor for the convergence speed of the EM algorithm. We differentiate between two cases based on the hyperparameter $R\in\{0,1\}$. For $R=1$, we sample $\boldsymbol{\pi}$ from the prior distribution $P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$ as starting point (re-initialization). For $R=0$, we use the last $\boldsymbol{\pi}$ as starting point (no re-initialization). The choice of $R$ has a significant impact on the convergence speed of the EM algorithm. We provide empirical results in the next section.
Algorithm 3 Latent Dirichlet Sampling
1:$(D_{1},\dots,D_{K})^{\top},\boldsymbol{\beta},\boldsymbol{\omega},B,\boldsymbol{y}$
2:$\Delta,\tau,$ $R$triangleright Hyperparameters
3:$B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$
4:$D\leftarrow\sum_{k=1}^{K}D_{k}$
5:$T\leftarrow\left\lceil D/B\right\rceil$
6:$B_{k}^{(t)}\leftarrow 0$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$
7:
8:$\boldsymbol{\alpha}\leftarrow(\frac{D_{1}}{D},\frac{D_{2}}{D},\dots,\frac{D_{K}}{D})^{\top}$
9:for $k=1$ to $K$ do
10:     $\alpha_{k}\leftarrow\alpha_{k}\cdot\exp\{\Delta\cdot\frac{(\omega_{k}-\bar{\boldsymbol{\omega}})}{s_{\boldsymbol{\omega}}}\}$
11:end for
12:$\boldsymbol{\pi}\sim$ Dir$(\boldsymbol{\alpha})$
13:$\boldsymbol{\pi}\leftarrow$ EmMap$(\boldsymbol{y},\boldsymbol{\pi},\boldsymbol{\beta},\boldsymbol{\alpha},\tau)$
14:
15:for each training step $t=1$ to $T$ do
16:     for $i=1$ to $B$ do
17:         $z_{i}\sim\text{Categorical}(\boldsymbol{\pi})$
18:         $B_{z_{i}}^{(t)}\leftarrow B_{z_{i}}^{(t)}+1$
19:         if $\sum_{t^{\prime}=1}^{t}B_{z_{i}}^{(t^{\prime})}=D_{z_{i}}$ then
20:              $\boldsymbol{\beta},\boldsymbol{\pi},\boldsymbol{\alpha}\leftarrow$ RemoveComponent$(z_{i},\boldsymbol{\beta},\boldsymbol{\pi},\boldsymbol{\alpha})$
21:
22:              if $R=1$ then
23:                  $\boldsymbol{\pi}\sim\text{Dir}(\boldsymbol{\alpha})$
24:              end if
25:
26:              $\boldsymbol{\pi}\leftarrow$ EmMap$(\boldsymbol{y},\boldsymbol{\pi},\boldsymbol{\beta},\boldsymbol{\alpha},\tau)$
27:         end if
28:     end for
29:end for
30:return $B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$
Figure 3: Sampling time for UGS and LDS with different $\Delta$.
V Simulation Results
This section provides a comprehensive evaluation of the proposed UGS and LDS methods in terms of model performance and training time. We compare PSL extended by our sampling methods with several existing frameworks, focusing on metrics such as test accuracy, convergence behavior, and training time per epoch (TPE). Furthermore, we investigate the effectiveness of LDS for mitigating stragglers, and analyze its time complexity under various conditions.
V-A Model Performance and Training Time
We trained a ResNet18 [35] model on the CIFAR10 [36] dataset using PSL with the proposed UGS and LDS (with $\Delta=0.0$) in the absence of stragglers. We compared the test accuracy with PSL using fixed and proportional local batch sizes like in [19]. For brevity we refer to this default sampling method as fixed proportional local sampling (FPLS). If the local batch sizes are fixed and identical like in [24], we refer to this sampling method as fixed local sampling (FLS). To ensure that all local batch sizes are integer, we round them to the nearest integer. If the nearest integer is 00, we set the batch size to $1$ to ensure that each client contributes to the training process. We also include other frameworks, namely SL, FL, and SFL, as well as CL as a baseline. The hyperparameter choices, hardware setup, and training details can be found in Appendix A.
We split the datasets either in an IID or non-IID manner, while considering different numbers of clients ${K\in\{16,32,64,128\}}$. To achieve a non-IID split, a Dirichlet-based data partition strategy is used, similar to the extended Dirichlet strategy in [15]. The strategy is applied to the datasets so that each client has exactly $2$ classes with strongly varying dataset sizes, simulating strong heterogeneity among the clients. The resulting client datasets are shown exemplary in Figure 4 for $K=16$ clients. We trained the model for 100 epochs with a global batch size of $B=128$ and report the maximum test accuracy achieved during training in Table II. Accuracies are averaged over $5$ runs with different random seeds and given in percentages. The best accuracy for each split and $K$ is highlighted in boldface. The non-IID split is denoted as $(\overline{\text{IID}})$ in the table. CL achieved a maximum accuracy of 90.25%. In Figure 5, we show the test accuracy per epoch for a run with $K=128$.
Figure 4: Example of a Non-IID setting for $K=16$ clients.
TABLE II: Test Accuracies for Different Frameworks
Test Acc. PSL (UGS) PSL (LDS) PSL (FPLS) FL SFL SSL
$K\downarrow$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$
16 90.43 90.24 89.85 90.80 89.89 87.14 87.65 44.20 88.10 48.04 90.18 21.26
32 90.75 89.98 90.66 89.90 90.05 74.52 86.31 49.84 86.96 49.71 90.68 33.18
64 90.11 90.17 90.21 90.72 90.26 71.75 84.33 48.65 84.77 53.87 90.53 30.92
128 90.10 89.95 90.37 90.28 90.07 67.32 81.96 48.34 81.80 47.59 88.06 40.40
Figure 5: Test accuracy per epoch for different frameworks training ResNet18 on CIFAR10 in IID and Non-IID settings.
We observe that PSL with UGS and LDS significantly outperforms the other frameworks in terms of test accuracy. Compared to PSL with FPLS, UGS and LDS increase the test accuracy by up to $34.1\%$ in the non-IID setting. Our proposed methods achieved a test accuracy close to the central learning baseline in both IID and non-IID settings, in some cases even surpassing it. The other frameworks show a significant drop in test accuracy in the non-IID setting and strong fluctuations, suggesting that the model training is unstable. The fluctuations are particularly pronounced for PSL with FPLS. In contrast, PSL with UGS and LDS shows a stable convergence behavior. The accuracies achieved by UGS and LDS are closely aligned. The results suggest that UGS and LDS are effective in mitigating the challenges of large effective batch sizes and non-IID data in PSL.
To investigate the performance gap between our methods and PSL with fixed local batch sizes, we compared the batch deviation (as defined in Equation 1) of UGS with FPLS and FLS. We calculated the mean and standard deviation of the batch deviation in the global batch sequence generated by the different methods. We considered different values of $B$ and $K$ under IID and non-IID settings, as shown in Figure 6. For each batch size $B$ we selected 5 different values for $K$ that were evenly distributed in $\{20,\dots,B\}$. In the IID scenario, all sampling strategies exhibited nearly identical means and standard deviations. This is consistent with Theorem 1, since $\beta_{0,m}\approx\beta_{k,m}$ for all $k$ and $m$. Differences between UGS, FPLS, and FLS were minimal, with UGS maintaining a nearly constant mean and standard deviation in both IID and non-IID cases. However, FPLS and FLS showed a marked increase in mean and standard deviation in the non-IID scenario. The worst batch deviation in non-IID cases was observed with FLS, while FPLS showed a lower mean and standard deviation that decreased with higher $K$, though still considerably higher than UGS. For UGS, a larger batch size consistently reduced both mean and standard deviation in both IID and non-IID cases, aligning with the deviation bound in Lemma 1. In contrast, for FPLS and FLS, this reduction was observed only in the IID case. This suggests the dominance of the bias term in the deviation bound of Lemma 2 in the non-IID setting. Our findings suggests that sampling methods with fixed local batch sizes introduce greater batch deviation to the training process, explaining the performance gap between UGS and FPLS.
Figure 6: Mean and standard deviation of the batch deviation for UGS, FLS, and FPLS using different $K$ and $B$.
Figure 7: Mean and standard deviation of the batch deviation for UGS and LDS using different $\Delta$, $K$, and $B$.
V-B Evaluation of LDS for Straggler Mitigation
To evaluate the effectiveness of the LDS method for straggler mitigation, we introduced stragglers into the training process by defining a straggler probability $p_{s}$, a minimum straggler delay $\omega_{\text{min}}\in\mathbb{N}$, and a maximum straggler delay $\omega_{\text{max}}\in\mathbb{N}$ with $\omega_{\text{min}}<\omega_{\text{max}}$. The delays are measured in milliseconds. We randomly selected a subset of clients to be stragglers with probability $p_{s}$ and introduced a delay $\boldsymbol{\omega}$ uniformly distributed between $\omega_{\text{min}}$ and $\omega_{\text{max}}$. For non-stragglers the delay time was set to zero. Each client $k$ waited for $\omega_{k}$ milliseconds before sending its data to the server. We trained the model again on the CIFAR10 dataset with the same hyperparameters using PSL with LDS. We measured the TPE for different straggler probabilities $p_{s}\in\{0.1,0.2,0.3\}$ and $\Delta\in\{0.5,1.0,1.5\}$ under IID and non-IID dataset splits. The results are shown in Table IV. TPEs are averaged over the $5$ runs and given in seconds. We included the TPE for $\Delta=0.0$ in the absence of stragglers as a baseline. The best TPE for each split, $K$, and $p_{s}$ is highlighted in boldface.
For $p_{s}=0.1$, the TPE increased drastically, especially for fewer clients $K$. Increasing $p_{s}$ further only slightly increased the TPE. Higher $\Delta$ values significantly reduced the TPE by up to $62\%$. For $p_{s}=0.1$ and $\Delta=1.5$ the TPE was only slightly higher than without stragglers. Higher $p_{s}$ made LDS less effective in reducing the TPE due to the increased straggler concentration in the global batch sequence, especially with fewer clients $K$. Figure 8 shows the processing times per batch and the total processing time for different $p_{s}$ and $\Delta$ with $K=128$ clients. Higher $\Delta$ led to faster decreases in the processing time. Higher $p_{s}$ reduced the rate of this decrease, setting a lower bound beyond which $\Delta$ had no effect.
We compared the test accuracy for each setting, as shown in Table III. Accuracies are averaged over $5$ runs and given in percentages. The best accuracy for each split, $K$ and $p_{s}$ is highlighted in boldface. The test accuracy remained close to the UGS baseline across all $p_{s}$ and $\Delta$, demonstrating robustness.
We compared the batch deviation of UGS and LDS in the same settings as before, using different values for $\Delta$, $K$, and $B$ under IID and non-IID settings. The results are shown in Figure 7. The difference in batch deviation was more pronounced than the difference in accuracy. In the IID case, the batch deviation was nearly identical for all cases. In the non-IID case, LDS with $\Delta=0.0$ closely aligned with UGS. Higher $\Delta$ increased mean and standard deviation, but significantly less than FLS and FPLS. Mean and standard deviation decreased with higher $K$ and $B$. The negligible effect on the model accuracy likely results from the low increase in batch deviation, which is not sufficient to impact the model training.
The significant TPE reduction without impacting accuracy, especially under a strong non-IID distribution, suggests that LDS is effective for straggler mitigation.
TABLE III: Test Accuracies for PSL with LDS
Test Acc. $p_{s}=0.1$ $p_{s}=0.2$ $p_{s}=0.3$
$\Delta=0.5$ $\Delta=1.0$ $\Delta=1.5$ $\Delta=0.5$ $\Delta=1.0$ $\Delta=1.5$ $\Delta=0.5$ $\Delta=1.0$ $\Delta=1.5$
$K\downarrow$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$ IID $\overline{\text{IID}}$
16 90.27 90.22 90.28 90.89 90.05 90.78 90.34 90.45 90.45 90.85 90.32 90.74 90.23 89.58 90.24 89.19 90.60 87.06
32 90.55 90.16 90.22 91.02 89.90 90.46 90.35 90.48 90.38 90.73 90.41 90.57 90.11 90.37 90.19 90.65 90.06 90.50
64 90.18 90.21 90.28 90.47 90.30 90.55 90.33 90.43 90.68 90.81 90.39 90.74 90.11 90.36 90.44 90.46 89.98 90.41
128 90.24 90.35 90.17 90.57 90.26 90.99 90.22 90.50 90.54 90.35 90.15 90.43 89.99 90.46 90.62 90.50 90.10 90.89
TABLE IV: Roundtimes for PSL with LDS
Roundtime $p_{s}=0.0$ $p_{s}=0.1$ $p_{s}=0.2$ $p_{s}=0.3$
$K\downarrow/\Delta\rightarrow$ $0.0$ $0.0$ $0.5$ $1.0$ $1.5$ $0.0$ $0.5$ $1.0$ $1.5$ $0.0$ $0.5$ $1.0$ $1.5$
16 9.71 27.13 14.95 12.99 11.94 28.09 19.95 14.82 12.98 33.82 22.33 20.23 18.23
32 11.57 35.93 16.05 14.03 13.68 38.47 23.13 19.20 17.09 39.07 26.24 23.61 19.07
64 13.62 38.29 21.83 18.63 16.19 41.26 26.64 21.07 19.51 39.53 28.56 23.95 23.15
128 15.68 41.55 24.70 20.45 18.11 43.82 28.83 23.94 21.92 44.11 32.26 26.30 24.26
Figure 8: Per-batch processing time and total time (i.e., TPE) in seconds for different $p_{s}$ and $\Delta$ with $K=128$ clients.
V-C Time Complexity of LDS
We measured the total number of EM iterations invoked during LDS for various values of $\Delta$, $K$, $N$, and $R$. We present our results for $p_{s}=0.1$ and $p_{s}=0.2$, respectively, as a 3D heatmap in Figure 2. We observe that reinitialization ($R=1$) significantly impacts the number of iterations. Without reinitialization ($R=0$), the EM algorithm requires on average less than a third of the iterations to converge. This indicates that the $\boldsymbol{\pi}$ from the previous iteration provides a good starting point for the next iteration. The highest number of iterations occurs when using $R=1$ with a high value of $\Delta$ and a small sample size. The EM algorithm needs more iterations to converge when the $\boldsymbol{\alpha}$ are adjusted more aggressively. In particular, with a small sample size. Poor initial values due to reinitialization strongly exacerbates this effect. For a higher number of clients $K$, the number of iterations increased in all cases. This is primarily because the EM algorithm must be re-run each time a client dataset depletes. The results show that for $R=0$ the choice of sample size $N$ has no significant impact on the number of iterations. This is due to the speedup achieved by the reformulation of the E-step in Section IV-C.
Lastly, we measured the total sampling time required for LDS and compared it with UGS. We used a straggler probability of $p_{s}=0.2$, a batch size of $B=128$, and different values of $\Delta$. The results, shown in Figure 3, indicate that the sampling time for LDS increases as $K$ increases. This is because the number of EM invokations and the number of computations for the $\hat{\boldsymbol{\gamma}}_{k}$ scales with $K$. However, LDS is still only slightly slower than UGS, even for high values of $\Delta$. Thus, LDS not only provides a significant reduction in TPE but also maintains a low computational overhead compared to UGS.
V-D Limitations
We outline some limitations of our methods that were observed during our analysis and simulations. First, we assume that the delay times do not change significantly during an epoch. However, this assumption may not hold in practice. The straggler mitigation may not be effective if the delay times change frequently. Furthermore, the inherent reliance on gradient averaging causes increased latency and communication overhead. Gradient averaging exacerbates the straggler effect, as the server must wait for all contributing clients to send their gradients before proceeding. Lastly, we observed that when the number of clients is at least as large as the global batch size, the local batch sizes can become very small, which leads to increased latency due to idle clients. The integration of UGS and LDS with other advanced techniques like local loss function design and client clustering could further enhance the performance of PSL, by removing the need for gradient averaging and improving the scalability of the framework. Other venues for future research include reducing the communication overhead and decreasing the training latency of PSL.
VI Conclusion
In this paper, we proposed the UGS and LDS methods as an extension to the PSL framework. Our methods addressed key challenges in PSL, including large effective batch sizes, non-IID data distributions, and the straggler effect. UGS mitigates the large effective batch size problem by allowing variable local batch sizes, decoupling the number of clients from the effective batch size. Our analysis demonstrated that UGS effectively reduces batch deviations, even under non-IID data settings. LDS is a generalization of UGS and balances the trade-off between batch deviation and training time using a single hyperparameter $\Delta$ that controls the trade-off. Our extensive simulations on the CIFAR10 dataset showed that PSL with UGS and LDS outperforms existing frameworks in terms of test accuracy and convergence stability. Specifically, our methods achieved a test accuracy close to the central learning baseline in the non-IID setting, while other frameworks experienced significant drops in performance. Compared to PSL with fixed local batch sizes, UGS and LDS increased the test accuracy by up to $34.1\%$ in the non-IID setting. The evaluation of LDS for straggler mitigation revealed substantial reduction in training time per epoch, achieving a reduction of up to $62\%$. Our findings suggest that UGS and LDS are promising solutions for enhancing PSL performance in real-world applications.
VII Acknowledgment
This work was supported by the Federal Ministry of Education and Research (BMBF, Germany) as part of NeuroSys: Impulse durch Anwendungen (Projekt D) - A under Grant 03ZU1106DA.
Appendix A Hyperparameters and Training Details
[…]
$L(\mathcal{B}^{(t)};\boldsymbol{w}_{s})=-\frac{1}{B}\sum_{i=1}^{B}\sum_{m=1}^{M}y_{i,m}\log\left(\hat{y}_{i,m}\right)$
where $\text{where }\hat{y}_{i}=f(f(\boldsymbol{x}_{i};\boldsymbol{w}_{c});\boldsymbol{w}_{s})$
[…]
Appendix B Proofs
Proof of Lemma 1
We have ${\mathbb{E}[Y_{m}]=B\beta_{0,m}}$. It holds that ${\mathbb{E}[\frac{Y_{m}}{B}]=\beta_{0,m}}$ and ${\text{Var}(\frac{Y_{m}}{B})=\frac{1}{B^{2}}\text{Var}(Y_{m})}$. The Lemma follows from the Chebychev inequality.
Proof of Lemma 2
According to the Markov inequality it holds that
$\mathbb{P}\left(\left(\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right)^{2}\geq\epsilon^{2}\right)\leq\frac{\mathbb{E}\left[\left(Y^{\prime}_{m}-B\beta_{0,m}\right)^{2}\right]}{B^{2}\epsilon^{2}}$
$\mathbb{E}$ $\left[\left(Y^{\prime}_{m}-B\beta_{0,m}\right)^{2}\right]=\mathbb{E}\left[\left(Y^{\prime}_{m}-\mathbb{E}\left[Y^{\prime}_{m}\right]+\mathbb{E}\left[Y^{\prime}_{m}\right]-B\beta_{0,m}\right)^{2}\right]$
$=\mathbb{E}\left[\left(Y^{\prime}_{m}-\mathbb{E}\left[Y^{\prime}_{m}\right]\right)^{2}\right]+2(\mathbb{E}\left[Y^{\prime}_{m}\right]-B\beta_{0,m})\mathbb{E}\left[Y^{\prime}_{m}-\mathbb{E}\left[Y^{\prime}_{m}\right]\right]$
$\quad+\left(\mathbb{E}\left[Y^{\prime}_{m}\right]-B\beta_{0,m}\right)^{2}$
$=\text{Var}(Y^{\prime}_{m})+\left(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}]\right)^{2},$
where the last equality follows from $\mathbb{E}\left[Y^{\prime}_{m}-\mathbb{E}\left[Y^{\prime}_{m}\right]\right]=0$ and $\mathbb{E}\left[Y_{m}\right]=B\beta_{0,m}$. Recognize that:
$\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\implies\left(\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right)^{2}\geq\epsilon^{2}.$
Therefore, the bound is:
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\text{Var}(Y^{\prime}_{m})+\left(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}]\right)^{2}}{B^{2}\epsilon^{2}}.$
Proof of Theorem 1
We first show that if either $\beta_{k,m}=\beta_{0,m}$ or $B_{k}=B\frac{D_{k}}{D}$ for all $k\in\mathcal{K}$ then $\mathbb{E}[Y^{\prime}_{m}]=\mathbb{E}[Y_{m}]$ for all $m\in\mathcal{M}$.
Case 1: $\beta_{k,m}=\beta_{0,m}:$
$\mathbb{E}[Y^{\prime}_{m}]$ $=\sum_{k\in\mathcal{K}}B_{k}\beta_{k,m}=\sum_{k\in\mathcal{K}}B_{k}\beta_{0,m}=\beta_{0,m}\sum_{k\in\mathcal{K}}B_{k}$
$=\beta_{0,m}B=\mathbb{E}[Y_{m}].$
Case 2: $B_{k}=B\frac{D_{k}}{D}:$
$D\beta_{0,m}=\sum_{k\in\mathcal{K}}D_{k}\beta_{k,m}$
$\iff$ $B\beta_{0,m}=\sum_{k\in\mathcal{K}}B\frac{D_{k}}{D}\beta_{k,m}=\sum_{k\in\mathcal{K}}B_{k}\beta_{k,m}$
$\iff$ $\mathbb{E}[Y^{\prime}_{m}]=\mathbb{E}[Y_{m}].$
Consequently, the bias term in Lemma 2 is zero and it holds that:
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y^{\prime}_{m})}{B^{2}\epsilon^{2}}.$
We now show that $\text{Var}(Y^{\prime}_{m})\leq\text{Var}(Y_{m})$ always holds. We have:
$\text{Var}(Y_{m})$ $=B\beta_{0,m}(1-\beta_{0,m})$
$=B\left(\sum_{k\in\mathcal{K}}\frac{B_{k}}{B}\beta_{k,m}\right)\left(1-\sum_{k\in\mathcal{K}}\frac{B_{k}}{B}\beta_{k,m}\right)$
$\geq B\sum_{k\in\mathcal{K}}\frac{B_{k}}{B}\beta_{k,m}(1-\beta_{k,m})$ (6)
$=\sum_{k\in\mathcal{K}}B_{k}\beta_{k,m}(1-\beta_{k,m})=\text{Var}(Y^{\prime}_{m})$
where Inequation (6) follows from Jensen’s inequality with the concave function $\varphi(x)=x(1-x)$. Equality holds if and only if ${\beta_{k,m}=\beta_{0,m}}$ for all $k\in\mathcal{K}$. Thus, it holds that
$\frac{\textnormal{Var}(Y^{\prime}_{m})}{B^{2}\epsilon^{2}}\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}$
which completes the proof.
Proof of Corollary 1
$\mathbb{E}[Y^{\prime}_{m}]=\sum_{k=1}^{K}B\pi_{k}\beta_{k,m}=\sum_{k=1}^{K}B_{k}\beta_{k,m}$
Where $B_{k}=B\pi_{k}=B\frac{D_{k}}{D}$. According to Theorem 1 the bound holds.
Proof of Proposition 1
We introduce a Lagrange multiplier $\lambda$ to ensure constraint $\sum_{k\in\mathcal{K}}\pi_{k}=1$, which yields the following objective function:
$\mathcal{L}(\boldsymbol{\pi},\lambda)=\mathcal{Q}(\boldsymbol{\pi},\boldsymbol{\pi}^{\text{old}})+\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})+\lambda\left(\sum_{k\in\mathcal{K}}\pi_{k}-1\right).$
Taking the derivative with respect to $\pi_{k}$ and setting it to zero, we get:
$\frac{\partial\mathcal{L}(\boldsymbol{\pi},\lambda)}{\partial\pi_{k}}\overset{!}{=}0$
$\iff$ $\sum_{n=1}^{N}\frac{\gamma_{n,k}}{\pi_{k}}+\frac{\alpha_{k}-1}{\pi_{k}}+\lambda=0$
$\iff$ $\sum_{n=1}^{N}\gamma_{n,k}+\alpha_{k}-1+\lambda\pi_{k}=0$ (7)
$\iff$ $\sum_{k^{\prime}=1}^{K}\sum_{n=1}^{N}\gamma_{n,k^{\prime}}+\sum_{k^{\prime}=1}^{K}(\alpha_{k^{\prime}}-1)+\sum_{k^{\prime}=1}^{K}\lambda\pi_{k^{\prime}}=0$
$\iff$ $N+\alpha_{0}-K+\lambda\sum_{k=1}^{K}\pi_{k}=0$
$\iff$ $\lambda=-N-\alpha_{0}+K.$
Inserting the result back into Equation 7 yields:
$\sum_{n=1}^{N}\gamma_{n,k}+\alpha_{k}-1+\lambda\pi_{k}=0$
$\iff$ $N_{k}+\alpha_{k}-1=(N+\alpha_{0}-K)\pi_{k}$
$\iff$ $\pi_{k}=\frac{N_{k}+\alpha_{k}-1}{N+\alpha_{0}-K}.$